what is the solution for this - log8 + log4
Found 3 solutions by Boreal, math_helper, greenestamps:
Answer by Boreal(15235)

This is log 4/log 8 or log (1/2) which is -log 2 or -0.3010
Minus logs are divided and logs with same bases may be combined. log(1/2) is log 1-log 2, but log 1 is 0, so it is - log 2.

Answer by math_helper(2461)
Answer by greenestamps(13237)



(1) rewrite the expression as log4 - log8

(2) Since logarithms are exponents, subtracting logarithms represents division. So log4 - log8 = log(4/8) = log(1/2)

Answer: -log8+log4 = log(1/2)

The other two tutors both gave "-log2" as the answer. The two answers are equivalent, so both are valid:

log(1/2) = log1-log2 = 0-log2 = -log2

or...

log(1/2) = log(2^-1) = -1(log2) = -log2